# Anas Anwarul Haq Khan1, Utkarsh Verma2, Prateek Chanda1, Ganesh Ramakrishnan1

### v1

@@ 1 Introduction
In this work, we define the problem as follows: given a video $v_{i}\in\{V_{1},V_{2},\cdots,V_{n}\}$, our goal is to use its title ($T_{v_{i}}$), transcript ($Tr_{v_{i}}$), and audio-derived annotations such as emotion recognition and speaker diarization ($A_{v_{i}}$) to identify and retain the most relevant content segments. We approach this task in a supervised setting, focusing on preserving semantic integrity and engagement. In contrast to prior work that centers solely on visual features [27, 28, 33, 1], our approach leverages pre-trained vision-language models (VLMs), with a strong emphasis on textual and audio-semantic signals.
@@ 2.1 Unsupervised Video Summarization
Unsupervised video summarization techniques [32, 3, 19, 6, 17, 29, 13, 18] do not rely on annotated summaries. Instead, they often employ heuristics such as clustering similar frames to reduce redundancy [10]. These methods are particularly effective in contexts like egocentric recordings and surveillance footage, where deterministic or handcrafted rules for frame selection can be established. While unsupervised models are appealing for their adaptability and reduced labeling cost, they tend to struggle with high-level semantic understanding and narrative coherence, especially in open-domain content.
@@ 2.2 Supervised Video Summarization
Supervised methods [31, 33, 7, 16, 9, 23, 15, 25, 30] learn from manually annotated video summaries [26, 11] to predict the importance of frames or segments. Early approaches like DSNet [33] rely on deep convolutional features and sequential modeling to exploit temporal dependencies, while VASNet [7] leverages self-attention for keyframe selection, offering better interpretability and lower complexity compared to recurrent networks. PGL-SUM [1] further refines this by integrating both local and global attention, improving the model’s ability to capture long-range dependencies.
More recent methods explore task-aware and multimodal designs. IPTNet [16] introduces a two-branch architecture that simultaneously handles summarization and moment localization via shared importance maps and cross-task propagation. CLIP-It [23] incorporates text guidance through multimodal transformers for generic and query-specific summarization. A2Summ [12] enhances multimodal alignment using contrastive losses and segment-aware embeddings, focusing on better synchronization between textual and visual cues.
@@ 3 Datasets
Dataset Num of Videos Num of Annotations Avg Duration Min Duration Max Duration SumMe [11] 25 15–18 2min 26s 32s 5min 24s TVSum [26] 50 20 3min 55s 1min 23s 10min 47s OVP [21] 50 5 1min 38s 46s 3min 29s YouTube [4] 39 5 3min 16s 9s 9min 32s
For our experiments, we focus on two widely-used video summarization benchmarks: SumMe [11] and TVSum [26]. These datasets are standard in the literature and provide well-annotated summaries with diverse content and evaluation protocols.
TVSum [26] consists of 50 videos sourced from YouTube, spanning 10 categories such as how-to tutorials, news, and documentaries, with 5 videos per category. Each video ranges from 1 to 5 minutes in duration. Frame-level importance scores are collected from 20 users per video by dividing the videos into uniform 2-second shots and aggregating user responses.
SumMe [11] contains 25 videos covering various real-life scenarios like holidays, cooking, and sports. The videos are typically between 1.5 to 6.5 minutes long. Human annotations are provided in the form of both interval-based and frame-level scores, with at least 15 user summaries per video.
We choose to evaluate exclusively on SumMe and TVSum due to their widespread use in the literature, the availability of multiple human-generated annotations, and the diversity of video content they represent. In addition to these two benchmarks, we reference several auxiliary datasets commonly used in prior works for augmentation and evaluation: YouTube [4] and Open Video Project (OVP) [21] datasets contain 39 and 50 videos respectively. YouTube videos are drawn from categories such as news, sports, and cartoons, while OVP contains a broader range including documentaries and educational content. These datasets provide additional training samples and are typically annotated with interval-based summaries. QFVS [8] provides ground-truth generic summaries for 4 long-form videos (3–5 hours each) from the UT Egocentric [19] dataset, captured from head-mounted cameras in uncontrolled environments. The summaries are created via user selection over temporally segmented shots, with final labels computed by averaging annotations from three users.
@@ 4.2 MuE-style Early Exit via Cosine-Similarity Confidence
We implement a dynamic early-exit strategy inspired by the MuE model [20] to accelerate inference in our vision-language summarization pipeline. The core idea is to bypass full-depth decoding for semantically simple inputs by enabling confident predictions at intermediate decoder stages. Unlike the original MuE, where intermediate exits are trained jointly, our approach applies early exit only during inference, with no modification to training procedures.
@@ 4.3 Audio Processing
  • Speaker Diarization: We employ the pyannote.audio toolkit [2] to detect and separate distinct speakers in the audio stream. This allows us to segment the video based on speaker turns and interactions, which is especially useful for analyzing dialogues and discussions.
  • Gender-Emotion Classification: Relying solely on transcripts may lead to the loss of affective information, particularly in videos with limited visual cues (e.g., voice-over narrations, commentary). To address this, we segment the audio into 3-second windows and use a deep learning model [^2] https://github.com/MiteshPuthran/Speech-Emotion-Analyzer trained on four publicly available datasets—SAVEE [14], RAVDESS [22], TESS [5], and CREMA-D [cao2014crema]—to classify both the speaker’s emotion and gender. This information enriches the representation of audio segments and provides a nuanced understanding of speaker affect and identity.
@@ 6.4 Knowledge Distillation and Multi-Stage Distillation for PaLI-Gemma-2 Models
  • VLM 28B: Acts as the teacher model; no distillation is applied, and the model is trained using standard cross-entropy loss $\mathcal{L}_{\text{CE}}(y,\hat{y})$.
@@ 6.5 Impact of Early Exit on Inference Time and Accuracy
[…]
@@ 8 Limitations
Random [24] 11.28
DSNet [33] 17.01
PGL-SUM [1] 17.80
VASNet [7] 19.30
IPTNet [16] 19.47
A2Summ [12] 23.96

### v2

# Anas Anwarul Haq Khan1, Utkarsh Verma2, Ganesh Ramakrishnan1

@@ 1 Introduction
In this work, we define the problem as follows: given a video $v_{i}\in\{V_{1},V_{2},\cdots,V_{n}\}$, our goal is to use its title ($T_{v_{i}}$), transcript ($Tr_{v_{i}}$), and audio-derived annotations such as emotion recognition and speaker diarization ($A_{v_{i}}$) to identify and retain the most relevant content segments. We approach this task in a supervised setting, focusing on preserving semantic integrity and engagement. In contrast to prior work that centers solely on visual features Zhang et al. (2016, 2018); Zhu et al. (2020); Apostolidis et al. (2021), our approach leverages pre-trained vision-language models (VLMs), with a strong emphasis on textual and audio-semantic signals.
@@ 2.1 Unsupervised Video Summarization
Unsupervised video summarization techniques Zhou et al. (2018); De Avila et al. (2011a); Lee et al. (2012); Elhamifar et al. (2015); Jung et al. (2019); Zhao and Xing (2014); He et al. (2019); Jung et al. (2020) do not rely on annotated summaries. Instead, they often employ heuristics such as clustering similar frames to reduce redundancy Gong and Liu (2000). These methods are particularly effective in contexts like egocentric recordings and surveillance footage, where deterministic or handcrafted rules for frame selection can be established. While unsupervised models are appealing for their adaptability and reduced labeling cost, they tend to struggle with high-level semantic understanding and narrative coherence, especially in open-domain content.
@@ 2.2 Supervised Video Summarization
Supervised methods Zhao et al. (2021); Zhu et al. (2020); Fajtl et al. (2019); Jiang and Mu (2022); Gong et al. (2014); Narasimhan et al. (2021); Ji et al. (2019); Sharghi et al. (2017); Zhao et al. (2018) learn from manually annotated video summaries Song et al. (2015); Gygli et al. (2014) to predict the importance of frames or segments. Early approaches like DSNet Zhu et al. (2020) rely on deep convolutional features and sequential modeling to exploit temporal dependencies, while VASNet Fajtl et al. (2019) leverages self-attention for keyframe selection, offering better interpretability and lower complexity compared to recurrent networks. PGL-SUM Apostolidis et al. (2021) further refines this by integrating both local and global attention, improving the model’s ability to capture long-range dependencies.
More recent methods explore task-aware and multimodal designs. IPTNet Jiang and Mu (2022) introduces a two-branch architecture that simultaneously handles summarization and moment localization via shared importance maps and cross-task propagation. CLIP-It Narasimhan et al. (2021) incorporates text guidance through multimodal transformers for generic and query-specific summarization. A2Summ He et al. (2023) enhances multimodal alignment using contrastive losses and segment-aware embeddings, focusing on better synchronization between textual and visual cues.
@@ 3 Datasets
Dataset Num of Videos Num of Annotations Avg Duration Min Duration Max Duration SumMe Gygli et al. (2014) 25 15–18 2min 26s 32s 5min 24s TVSum Song et al. (2015) 50 20 3min 55s 1min 23s 10min 47s OVP Li et al. (2010) 50 5 1min 38s 46s 3min 29s YouTube De Avila et al. (2011b) 39 5 3min 16s 9s 9min 32s
For our experiments, we focus on two widely-used video summarization benchmarks: SumMe Gygli et al. (2014) and TVSum Song et al. (2015). These datasets are standard in the literature and provide well-annotated summaries with diverse content and evaluation protocols.
TVSum Song et al. (2015) consists of 50 videos sourced from YouTube, spanning 10 categories such as how-to tutorials, news, and documentaries, with 5 videos per category. Each video ranges from 1 to 5 minutes in duration. Frame-level importance scores are collected from 20 users per video by dividing the videos into uniform 2-second shots and aggregating user responses.
SumMe Gygli et al. (2014) contains 25 videos covering various real-life scenarios like holidays, cooking, and sports. The videos are typically between 1.5 to 6.5 minutes long. Human annotations are provided in the form of both interval-based and frame-level scores, with at least 15 user summaries per video.
We choose to evaluate exclusively on SumMe and TVSum due to their widespread use in the literature, the availability of multiple human-generated annotations, and the diversity of video content they represent. In addition to these two benchmarks, we reference several auxiliary datasets commonly used in prior works for augmentation and evaluation: YouTube De Avila et al. (2011b) and Open Video Project (OVP) Li et al. (2010) datasets contain 39 and 50 videos respectively. YouTube videos are drawn from categories such as news, sports, and cartoons, while OVP contains a broader range including documentaries and educational content. These datasets provide additional training samples and are typically annotated with interval-based summaries. QFVS Fu et al. (2020) provides ground-truth generic summaries for 4 long-form videos (3–5 hours each) from the UT Egocentric Lee et al. (2012) dataset, captured from head-mounted cameras in uncontrolled environments. The summaries are created via user selection over temporally segmented shots, with final labels computed by averaging annotations from three users.
@@ 4.2 MuE-style Early Exit via Cosine-Similarity Confidence
We implement a dynamic early-exit strategy inspired by the MuE model Li et al. (2023) to accelerate inference in our vision-language summarization pipeline. The core idea is to bypass full-depth decoding for semantically simple inputs by enabling confident predictions at intermediate decoder stages. Unlike the original MuE, where intermediate exits are trained jointly, our approach applies early exit only during inference, with no modification to training procedures.
@@ 4.3 Audio Processing
  • Speaker Diarization: We employ the pyannote.audio toolkit Bredin et al. (2020) to detect and separate distinct speakers in the audio stream. This allows us to segment the video based on speaker turns and interactions, which is especially useful for analyzing dialogues and discussions.
  • Gender-Emotion Classification: Relying solely on transcripts may lead to the loss of affective information, particularly in videos with limited visual cues (e.g., voice-over narrations, commentary). To address this, we segment the audio into 3-second windows and use a deep learning model [^2] https://github.com/MiteshPuthran/Speech-Emotion-Analyzer trained on four publicly available datasets—SAVEE Jackson and Haq (2014), RAVDESS Livingstone and Russo (2018), TESS Dupuis and Pichora-Fuller (2010), and CREMA-D cao2014crema—to classify both the speaker’s emotion and gender. This information enriches the representation of audio segments and provides a nuanced understanding of speaker affect and identity.
@@ 6.4 Knowledge Distillation and Multi-Stage Distillation for PaLI-Gemma-2 Models
[…]
@@ 6.5 Impact of Early Exit on Inference Time and Accuracy
Early Exit achieves a  $\approx\textbf{~21\% reduction}$ in inference time, demonstrating its effectiveness in accelerating model response. However, this comes at the cost of a  2.1% absolute drop in F1 score, which is a significant degradation compared to other configurations. We hypothesize that this drop stems from architectural limitations of the 3B model, which was not extensively optimized for early exit during inference. Further architecture-aware tuning may mitigate this performance loss in future.
@@ 8 Limitations
Random Otani et al. (2019) 11.28
DSNet Zhu et al. (2020) 17.01
PGL-SUM Apostolidis et al. (2021) 17.80
VASNet Fajtl et al. (2019) 19.30
IPTNet Jiang and Mu (2022) 19.47
A2Summ He et al. (2023) 23.96